# Curing Semantic Drift: A Dynamic Approach to Grounding Generation in Large Vision-Language Models

### v3

@@ Abstract
Large Vision-Language Models (LVLMs) face a tug-of-war between powerful linguistic priors and visual evidence, often leading to “semantic drift”—the progressive detachment from visual input that we identify as the root cause of hallucination. While several existing training-free decoding strategies have achieved considerable success, they still suffer from inherent limitations. Many are computationally prohibitive, requiring multiple forward passes through the entire LVLM, while others rely on indirect, heuristic-based proxies that are unreliable correlates for a direct semantic conflict. We propose Dynamic Logits Calibration (DLC), a novel training-free framework that is the first to cure semantic drift in a direct, dynamic, and efficient manner. At each decoding step, DLC introduces a real-time visual referee that performs a dual-aspect visual alignment check: (1) it assesses the intrinsic visual relevance of a candidate token and (2) its contextual visual coherence. By dynamically balancing these two checks and evaluating them against an adaptive baseline, DLC surgically modulates the output logits to favor grounded tokens. Extensive experiments show DLC significantly outperforms existing methods in mitigating hallucinations while, crucially, maintaining high inference efficiency by avoiding costly multiple LVLM forward passes. Our work presents a powerful and practical solution for building more reliable and visually-grounded LVLMs. Code will be released on https://github.com/JiaheChen2002/DLC.
@@ 1 Introduction
Figure 1: Comparison of decoding methods across three dimensions: Speed, Correctness, and Detailedness using LLaVA-1.5 [llava], with all experiments conducted on a single NVIDIA A100 GPU. The vertical axis represents correctness, while the horizontal axis shows tokens per second (TPS). Bubble size corresponds to Detailedness of responses.
@@
In recent years, LVLMs [shikra, InstructBLIP, llava-next, Blip-2, llava, Qwen2-VL, mplug-owl2, minigpt4] have revolutionized multimodal understanding [Qwen-VL, survey, lisa, liu2024improved] by integrating powerful visual encoders [clip, siglip, fgclip] with Large Language Models (LLMs) [qwen, vicuna, llama, llama2]. However, the impressive autoregressive generation capabilities of these models conceal an inherent conflict: a fundamental tension between following the strong statistical patterns learned from vast text corpora (linguistic priors) and remaining faithful to the visual evidence in the input image [survey, perturbollava, opera, vcd, vasparse]. This conflict gives rise to a phenomenon we term semantic drift: during the course of generating long text, the model’s output gradually drifts away from the visual context, eventually succumbing to the inertia of its linguistic patterns. This drift is the primary driver of hallucination, the generation of plausible but factually ungrounded text, which remains a critical bottleneck for deploying LVLMs in high-stakes domains like medical diagnosis [meddetecting, llavamed] and autonomous systems [driving, reason2drive].
To combat hallucination, the community has explored various training-free decoding strategies, but their approaches to curing the underlying semantic drift fall into flawed categories. Some attempt intervention by only contrasting internal decoder states (e.g., DoLa [dola], DeCo [deco]), ignoring the external visual input entirely and thus failing to address the visual-textual drift. Others track indirect, heuristic-based proxies for drift, such as attention patterns [opera] or latent-space feature spectrums [iat, latenspacesteer], which are unreliable correlates for a direct semantic conflict. The most common category, contrastive-input methods (e.g., VCD [vcd], ICD [icd], SID [sid]), require computationally prohibitive multiple forward passes through the entire LVLM. Finally, post-hoc correction methods [amoh, woodpecker, degf] are inefficient as they only clean up the error after semantic drift has already occurred.
We argue the gap in the literature is the lack of a method that can cure semantic drift in a way that is simultaneously direct, dynamic, efficient, and performs a sufficiently fine-grained semantic check. While some works (e.g., CGD [gcd]) validate using CLIP for direct semantic guidance, their intervention is coarse-grained (at the sentence-level), missing the dual-aspect nature of token-level failure. Therefore, to pinpoint the direct, token-level signal of this drift, our fine-grained visual analysis (Figure 2) reveals a critical failure point: at the moment of hallucination, the model confidently selects a visually ungrounded token, despite the availability of more faithful alternatives. This divergence marks the precise moment of semantic drift—where linguistic priors hijack the generation process away from the visual evidence. This motivates our central research question: can we design a mechanism to intervene in real-time, at the token-level, by directly measuring visual fidelity at these critical junctures?
Motivated by this insight, we propose Dynamic Logits Calibration (DLC), a novel, efficient, and training-free decoding framework designed to resolve semantic drift at its source. The core idea of DLC is to introduce a real-time visual referee that, before generating each token, performs our novel dual-aspect visual alignment check on each candidate. Specifically, it assesses both: 1) the candidate’s intrinsic visual relevance and 2) its contextual visual coherence. These two scores are then combined and evaluated against a dynamic historical baseline to compute the Relative Visual Advantage (RVA). Through an adaptive weighting mechanism, DLC integrates this RVA directly into the original logits, dynamically boosting the probability of visually-grounded tokens. This mechanism allows DLC to precisely suppress hallucinations driven by linguistic priors without compromising fluency and, crucially, without any costly additional forward passes, thus preserving high inference efficiency. As a compelling preview of our results, Figure 1 illustrates that DLC is unique among decoding strategies, simultaneously achieving state-of-the-art correctness and superior inference speed, all while generating detailed and rich descriptions.
@@
  • We propose DLC, a novel training-free decoding framework that introduces a dual-aspect visual grounding mechanism (intrinsic relevance and contextual coherence) to efficiently cure semantic drift at each token step.
  • We demonstrate through extensive experiments that DLC sets a new state-of-the-art in hallucination mitigation across various LVLMs and benchmarks, while maintaining superior inference speed, offering a practical and powerful path toward more trustworthy LVLMs.
@@ 2.1 The Challenge of Hallucination in LVLMs
Hallucination in LVLMs occurs when generated text contradicts the visual input, arising from a tension between the model’s internal knowledge and external visual evidence. While users expect responses to be strictly conditioned on the image, LVLMs often default to powerful linguistic priors acquired from their training corpus. This tendency causes the generative process to favor statistically likely but visually ungrounded phrases, leading to a progressive semantic drift away from the visual reality as the text sequence unfolds. Our work confronts this core challenge by proposing a mechanism to dynamically re-anchor the generation process to visual evidence at inference time (see Appendix 7.1 for additional background on LVLMs).
@@ 2.2 A Taxonomy of Mitigation Strategies
The prior approaches to mitigating hallucination broadly fall into three paradigms. 1) Data-driven optimization [perturbollava, hacl, rit, gfaif, rlhf_v, rlaif_v, lessismore, hadpo] employs techniques like reinforcement learning and instruction tuning, but these methods demand expensive annotated data and risk overfitting. 2) Post-hoc correction mechanisms [Pelican, LogicCheckGPT, woodpecker, degf, amoh] utilize auxiliary networks or external knowledge to rectify errors, though they introduce significant inference latency. 3) Decoding strategy refinement [halc, gcd, opera, sid, vcd, latenspacesteer, iat, icd, marine, vasparse] includes methods like VCD [vcd], which uses contrastive decoding to attenuate textual bias, and OPERA [opera], which employs a retrospection mechanism. Despite their varied mechanisms, these paradigms share a critical limitation: they apply static or pre-defined policies to what is fundamentally a dynamic problem. They lack the real-time, content-aware adaptiveness required to counteract the progressive semantic drift that causes hallucination. Our DLC is designed to fill this gap by introducing a truly dynamic calibration mechanism that directly confronts and corrects for semantic drift as it happens.
@@ 3 Preliminary and Motivation
Figure 2: Visualization of semantic drift and token selection in LVLMs. The central plot tracks CCTA against a Historical Baseline $\bar{B}_{t}$, contrasting visually faithful (green) and hallucination (red) phases. The surrounding bar charts provide snapshots of token selection.
@@ 3.3 Motivation
  • Historical Baseline ($\boldsymbol{\bar{B}_{t}}$): A smoothed average of recent CCTA scores, representing the established level of visual coherence the model has recently maintained.
@@
This setup, by directly comparing logits against CCTA and ITA, allows us to visualize the exact conflict between the model’s linguistic bias and the grounded visual evidence.
The Diagnosis: Linguistic Priors Hijacking Generation. Our analysis reveals a clear pattern of behavior. During visually-grounded generation (green phase in Fig 2), the CCTA trajectory remains high, closely tracking the historical baseline, indicating successful selection of visually coherent tokens. However, as the model enters a hallucinatory phase (red phase), a critical divergence occurs. The CCTA score drops sharply, visually confirming the semantic drift. More critically, the surrounding bar charts reveal the direct cause: the model, driven by its powerful linguistic priors, confidently assigns the highest logit score to a visually inappropriate token with low visual alignment (ITA and potential CCTA). This detrimental selection occurs even when other tokens with far superior visual fidelity are available, which the model disregards, blinded by its linguistic bias.
@@ 4.1 Real-time Visual Alignment Assessment
At each decoding step $t$, DLC first establishes a dynamic benchmark of visual consistency from the recent context, against which it then scores each candidate token.
Establishing a Dynamic Visual Baseline. To effectively detect semantic deviations, we first establish a dynamic baseline of the model’s typical visual coherence. We maintain a Historical Baseline ($\bar{B}_{t}$), which represents the model’s recently established standard of visual fidelity. This baseline is calculated as a smoothed average of the History Contextual Token Alignment (HCTA) scores, computed from a recent window of $N$ tokens. The HCTA score measures the CLIP similarity between the image $v$ and the immediately preceding text window $y_{t-N:t-1}$:
@@ 4.2 Adaptive Logit Modulation
A positive RVA indicates the token pulls the generation towards visual truth and vice versa. This normalized metric provides a clear, actionable signal for intervention.
@@ 5 Experiments
Table 1: CHAIR hallucination evaluation results. This table evaluates model hallucination using the CHAIRS and CHAIRI metrics, where lower values indicate reduced hallucination and improved performance. TPS indicates tokens per second during generation. Bold and underlined values indicate the best and second-best results, respectively.
Max
Token
Methods LLaVA-1.5 InstructBLIP MiniGPT-4
$C_{S}$ $\downarrow$ $C_{I}$ $\downarrow$ TPS $\uparrow$ $C_{S}$ $\downarrow$ $C_{I}$ $\downarrow$ TPS $\uparrow$ $C_{S}$ $\downarrow$ $C_{I}$ $\downarrow$ TPS $\uparrow$
64 Nucleus 25.2 9.11 17.4 31.3 11.3 21.9 25.5 9.72 20.0
VCD 24.9 8.15 8.92 27.8 9.56 10.7 22.5 8.27 10.2
ICD 24.0 8.12 8.83 29.3 9.63 9.92 22.8 8.24 9.78
SID 21.9 6.84 8.82 27.7 9.14 9.86 23.7 8.21 9.14
DLC(Ours) 20.9 6.22 16.1 25.4 8.38 13.1 22.3 7.84 13.1
512 Nucleus 52.3 16.4 17.4 56.2 17.6 21.1 35.2 12.2 19.4
VCD 56.5 17.0 8.74 62.0 18.1 10.4 31.9 10.4 10.0
ICD 52.1 15.4 8.21 65.4 19.7 10.1 33.2 10.5 9.56
SID 51.1 14.4 8.61 58.9 16.8 9.88 33.5 11.1 8.92
DLC(Ours) 38.4 10.8 14.7 51.8 15.2 12.7 31.6 10.0 10.8
@@ 5.1 Experimental Settings
Benchmarks. To evaluate our proposed method, we employed a suite of established benchmarks. These include: (1) CHAIR [chair] and (2) POPE [pope], for quantifying object hallucination in image captions; (3) SHR Evaluations [hadpo], for detecting more fine-grained hallucination types; (4) GPT-4o assisted evaluation [woodpecker], for a comprehensive analysis of hallucinations and text quality; (5) MME benchmark [mme], for assessing the LVLM’s general ability; (6) LLaVA-Bench-in-the-Wild [llava], for performing targeted case studies on complex tasks. Appendix 7.2 provides detailed descriptions of the benchmarks.
Baselines. To evaluate our training-free LVLM decoding strategies, we have compared them with standard decoding methods, including nucleus sampling and greedy sampling. We also assessed the performance of our method against several state-of-the-art decoding approaches: VCD [vcd], ICD [icd], SID [sid], and OPERA [opera]. For comprehensive comparisons, we implemented VCD, ICD, SID, and DLC using both nucleus sampling (Top-pp=1). The reported performance of these baselines is based on our re-implementation using their publicly available code.
@@ 5.2 Experimental Results
POPE Evaluation. To assess the generalizability of our method beyond open-ended descriptive generation, we extended DLC to the VQA domain. We evaluated its performance on the POPE benchmark [pope], which is designed to probe object existence hallucinations through yes/no questions. As presented in Table 2, DLC achieves state-of-the-art results among training-free methods across all three challenging settings: random, popular, and adversarial. This robust performance demonstrates the effectiveness and versatility of our calibration framework, confirming its applicability to more constrained, reasoning-based tasks.
@@
Random Popular Adversarial Random Popular Adversarial Random Popular Adversarial
@@
SHR Evaluation. Following [opera, sid, hadpo], we evaluate on 200 images from the VG dataset using the prompt “Please describe this image in detail” with a maximum of 512 new tokens generated. Additionally, we evaluated the model’s generation speed using tokens per second (TPS). As illustrated in Figure 4, we observe that our DLC achieves state-of-the-art results on metrics related to both hallucinated sentences and words. This indicates that DLC can effectively assist models in mitigating hallucination issues stemming from strong semantic priors. Consistent with observations for OPERA [opera], we also find that for certain models, DLC may reduce the length of the generated text, which can be attributed to the reduction of superfluous hallucinatory content.
@@
Figure 4: SHR evaluation results. Six aspects are analyzed, including the number of sentences per image (SPI), the tokens per second (TPS), the number of hallucinated sentences per image (HSPI), the number of hallucinated words per image (HWPI), the ratio of hallucinated sentences (HSR), and the ratio of hallucinated words (HWR). Larger radar indicates better performance.
GPT-4o Assisted Evaluation. To further analyze the hallucinations and text quality for open-ended generation tasks, we employ GPT-4o as an advanced evaluation proxy. For evaluation, models generate detailed descriptions for 500 randomly selected COCO [mscoco] images, prompted with “Please describe this image in detail.” Comparative analyses are performed, contrasting our DLC with several established baselines, including Nucleus Sampling, VCD [vcd], OPERA [opera], and SID [sid]. The empirical results, summarized in Table 3(a), indicate that DLC consistently outperforms these baselines on most metrics. In particular, DLC demonstrates a significant improvement in correctness over conventional sampling-based decoding, while maintaining a comparable degree of detail. The close alignment of GPT-4o’s perceptual and reasoning faculties with human judgment lends substantial credibility to these findings, suggesting that the observed enhancements in hallucination reduction and text quality are indicative of genuine improvements from a human evaluative perspective.
MME Evaluation. To further assess DLC’s impact on the general abilities of LVLMs, we evaluated its performance on the MME benchmark [mme]. The results on the LLaVA-1.5 model, averaged over five random seeds, demonstrate that DLC preserves the model’s performance on core perceptual grounding tasks, as illustrated in Figure 5. Full MME results are provided in the Appendix 7.6.
Figure 5: Evaluation on the subset of the MME benchmark.
Case Study on LLaVA-Bench. To complement our dedicated metrics, we also evaluated DLC on the challenging LLaVA-Bench-in-the-Wild dataset [llava]. We first present a comprehensive GPT-4o evaluation of all generation outputs, which quantitatively assesses correctness and detail. As detailed in Table 3(b), these results substantiate DLC’s superior performance over competing methods. Furthermore, we provide an in-depth qualitative analysis of case studies in Appendix 7.10. These cases visually demonstrate how the DLC approach markedly reduces hallucinatory content and enhances descriptive detail under identical image and prompt conditions.
Table 3: Assessing hallucination with GPT-4o on two benchmarks. Performance is assessed on correctness ($C$) and detailedness ($D$).
(a) Results on COCO [mscoco]

### v4

# Curing Semantic Drift: A Dynamic Approach to Grounding Generation in LVLMs

@@ Abstract
Large Vision-Language Models (LVLMs) face a tug-of-war between powerful linguistic priors and visual evidence, often leading to semantic drift: a progressive detachment from the input image that can abruptly emerge at specific decoding steps. Through a token-level diagnosis, we show that hallucination is frequently triggered not by the absence of grounded candidates, but by a failure of selection—the model chooses a linguistically convenient yet visually unfaithful token even when better grounded alternatives exist. Motivated by this insight, we propose Dynamic Logits Calibration (DLC), a training-free decoding framework that introduces a lightweight visual referee to intervene exactly when drift happens. At each step, DLC performs a dual-aspect grounding check on top-$k$ candidates: (1) it assesses the intrinsic visual relevance of a candidate token and (2) its contextual visual coherence. These signals are evaluated against an adaptive historical baseline to compute a relative visual advantage, which is then used to dynamically calibrate logits and favor grounded tokens. Extensive experiments on CHAIR, POPE, SHR, GPT-4o evaluation, and MME demonstrate that DLC consistently reduces hallucinations across multiple LVLMs while preserving response quality. Further analyses validate robustness to different vision backbones and demonstrate a favorable trade-off between output quality and computational cost as the candidate pool size varies.
@@ 1 Introduction
Figure 1: Semantic drift manifests as a detectable collapse of visual alignment and a failure of token selection. The central plot tracks the Current Contextual Token Alignment (CCTA) against an adaptive historical baseline $\bar{B}_{t}$, contrasting visually faithful (green) and hallucination (red) phases. The accompanying token-level snapshots compare top-candidate preferences with their visual alignment, revealing that hallucination is often triggered when a linguistically plausible but visually unfaithful token is selected over grounded alternatives.
@@
In recent years, LVLMs [shikra, InstructBLIP, llava-next, Blip-2, llava, Qwen2-VL, mplug-owl2, minigpt4] have revolutionized multimodal understanding [Qwen-VL, survey, lisa, liu2024improved] by integrating powerful visual encoders [clip, siglip, fgclip] with Large Language Models (LLMs) [qwen, vicuna, llama, llama2]. However, the impressive autoregressive generation capabilities of these models conceal an inherent conflict: a fundamental tension between following the strong statistical patterns learned from vast text corpora (linguistic priors) and remaining faithful to the visual evidence in the input image [survey, perturbollava, opera, vcd, vasparse]. This conflict gives rise to a phenomenon we term semantic drift: as decoding proceeds—especially for long-form generation—the model’s outputs progressively deviate from the visual context and become increasingly dominated by linguistic priors. This drift is a major driver of hallucination, the generation of plausible but factually ungrounded text, which remains a critical bottleneck for deploying LVLMs in high-stakes domains like medical diagnosis [meddetecting, llavamed] and autonomous systems [driving, reason2drive].
To cure this drift, we dissect its micro-dynamics and argue that hallucination is not a static error rate but a trajectory failure during decoding. As previewed in Figure 1, a token-level visualization of the Current Contextual Token Alignment (CCTA) reveals two key observations. First, drift has a moment: the alignment remains stable when generation is grounded, but collapses sharply at the onset of hallucination, suggesting a detectable transition rather than uniform noise. Second, drift is often a selection problem: around these critical steps, visually grounded tokens can still appear among top candidates, yet the model selects a linguistically convenient but visually unfaithful token. This indicates that the model is not necessarily blind, but can be overruled by linguistic priors during token choice. (A detailed analysis is provided in the Section 3.3.)
This diagnosis motivates a simple principle: if hallucination stems from choosing the wrong token at the right time, then we do not need to retrain the LVLM or perform multi-pass contrastive decoding; instead, we need a mechanism that injects direct visual grounding signals into the token selection process in real time. We propose Dynamic Logits Calibration (DLC), a lightweight, training-free decoding framework that acts as a visual referee. At each step, DLC evaluates candidate tokens from two complementary aspects—Intrinsic Relevance (is the concept visually present?) and Contextual Coherence (is it consistent with the current visual narrative?)—and dynamically calibrates logits to prevent linguistic priors from hijacking token choice, without disrupting the generation flow.
@@
  • We propose DLC, a novel training-free decoding framework that introduces a dual-aspect visual grounding mechanism (intrinsic relevance and contextual coherence) to effectively cure semantic drift at each token step.
  • We demonstrate through extensive experiments that DLC sets a new state-of-the-art in hallucination mitigation across various LVLMs and benchmarks, offering a practical path toward trustworthy LVLM decoding.
@@ 2.1 The Challenge of Hallucination in LVLMs
Hallucination in LVLMs occurs when generated text contradicts the visual input, arising from a tension between the model’s internal knowledge and external visual evidence. While users expect responses to be strictly conditioned on the image, LVLMs often default to powerful linguistic priors acquired from their training corpus. This tendency causes the generative process to favor statistically likely but visually ungrounded phrases, leading to a progressive semantic drift away from the visual reality as the text sequence unfolds. Our work confronts this core challenge by proposing a mechanism to dynamically re-anchor the generation process to visual evidence at inference time (see Appendix for additional background on LVLMs).
@@ 2.2 A Taxonomy of Mitigation Strategies
Prior approaches to mitigating hallucination in LVLMs can be broadly grouped by when and how they intervene in the generation pipeline. 1) Data-driven optimization [perturbollava, hacl, rit, gfaif, rlhf_v, rlaif_v, lessismore, hadpo] improves faithfulness by modifying model behavior via instruction tuning or RLHF. While effective, these methods typically require substantial curated data and careful alignment to avoid overfitting or regressions across tasks. 2) Post-hoc correction [Pelican, LogicCheckGPT, woodpecker, amoh] detects and rectifies errors via auxiliary verifiers, tool use, or external knowledge. Although this improves reliability, it functions reactively—correcting errors only after they occur—which increases system complexity and inference latency. 3) Decoding-time refinement offers a training-free alternative that adjusts generation on the fly. This line can be further categorized by the source of guidance:
  • Internal self-guided decoding. Approaches such as VCD [vcd], SID [sid], and OPERA [opera] exploit signals from the LVLM itself to discourage hallucination. While self-contained, they often require additional LVLM evaluations or specialized decoding procedures, and typically rely on heuristic signals rather than explicit semantic verification.
  • External-guided decoding. Methods like HALC [halc], MARINE [marine], DeGF [degf], and CGD [gcd] use external guidance to steer generation. Such guidance provides explicit semantic signals but often introduces extra modules or search overhead. Furthermore, many of these methods operate at a coarse granularity, lacking the agility to intervene precisely at the token level.
Despite their diversity, many existing methods rely on static penalties or incur prohibitive computation to address what is fundamentally a dynamic, token-level failure mode. DLC fills this gap by performing direct, token-level calibration with a lightweight visual referee: it injects grounded signals (intrinsic relevance and contextual coherence) into token selection exactly as semantic drift emerges. Crucially, DLC achieves this without the additional forward passes through the LVLM backbone required by multi-pass contrastive decoding baselines, ensuring both efficiency and precision.
@@ 3.3 Motivation
  • Historical Baseline ($\boldsymbol{\bar{B}_{t}}$): A smoothed average of recent CCTA scores, reflecting the model’s established visual coherence.
@@
Comparing logits against CCTA and ITA allows us to visualize the exact conflict between the model’s linguistic bias and the grounded visual evidence.
The Diagnosis: Linguistic Priors Hijacking Generation. Our analysis reveals a clear pattern of behavior. During visually-grounded generation (green phase in Fig 1), the CCTA trajectory remains high, closely tracking the historical baseline, indicating successful selection of visually coherent tokens. However, as the model enters a hallucinatory phase (red phase), a critical divergence occurs. The CCTA score drops sharply, visually confirming the semantic drift. More critically, the surrounding bar charts reveal the direct cause: the model, driven by its powerful linguistic priors, confidently assigns the highest logit score to a visually inappropriate token with low visual alignment. This detrimental selection occurs even when other tokens with far superior visual fidelity are available, which the model disregards, blinded by its linguistic bias.
@@ 4.1 Real-time Visual Alignment Assessment
At step $t$, DLC first establishes a dynamic benchmark of visual consistency from the recent context, against which it then scores each candidate token.
Establishing a Dynamic Visual Baseline. To effectively detect semantic deviations, we first establish a dynamic baseline of the model’s visual coherence. We maintain a Historical Baseline ($\bar{B}_{t}$), which represents the model’s recently established standard of visual fidelity. This baseline is calculated as a smoothed average of the History Contextual Token Alignment (HCTA) scores, computed from a recent window of $N$ tokens. The HCTA score measures the CLIP similarity between the image $v$ and the immediately preceding text window $y_{t-N:t-1}$:
@@ 4.2 Adaptive Logit Modulation
A positive RVA indicates the token pulls the generation towards visual truth and vice versa. This metric provides a clear, actionable signal for intervention.
@@ 5 Experiments
Table 1: CHAIR hallucination evaluation results. This table evaluates model hallucination using the CHAIRS and CHAIRI metrics, where lower values indicate reduced hallucination and improved performance. Bold and underlined values indicate the best and second-best results, respectively.
Max Token Methods LLaVA-1.5 InstructBLIP MiniGPT-4
64 Nucleus 25.2 9.11 31.3 11.3 25.5 9.72
VCD 24.9 8.15 27.8 9.56 22.5 8.27
ICD 24.0 8.12 29.3 9.63 22.8 8.24
SID 21.9 6.84 27.7 9.14 23.7 8.21
DLC(Ours) 20.9 6.22 25.4 8.38 22.3 7.84
512 Nucleus 52.3 16.4 56.2 17.6 35.2 12.2
VCD 56.5 17.0 62.0 18.1 31.9 10.4
ICD 52.1 15.4 65.4 19.7 33.2 10.5
SID 51.1 14.4 58.9 16.8 33.5 11.1
DLC(Ours) 38.4 10.8 51.8 15.2 31.6 10.0
@@ 5.1 Experimental Settings
Benchmarks. To evaluate our proposed method, we employed a suite of established benchmarks. These include: (1) CHAIR [chair] and (2) POPE [pope], for quantifying object hallucination in image captions; (3) SHR Evaluations [hadpo], for detecting more fine-grained hallucination types; (4) GPT-4o assisted evaluation [woodpecker], for a comprehensive analysis of hallucinations and text quality; (5) MME benchmark [mme], for assessing the LVLM’s general ability; (6) LLaVA-Bench-in-the-Wild [llava], for performing targeted case studies on complex tasks. Appendix provides detailed descriptions of the benchmarks.
Baselines. To evaluate our training-free LVLM decoding strategies, we have compared them with standard decoding methods and several state-of-the-art decoding approaches: VCD [vcd], ICD [icd], SID [sid], and OPERA [opera]. For comprehensive comparisons, we implemented VCD, ICD, SID, and DLC using both nucleus sampling (Top-pp=1). The reported performance of these baselines is based on our re-implementation using their publicly available code.
@@ 5.2 Experimental Results
POPE Evaluation. To assess the generalizability of our method beyond open-ended descriptive generation, we extended DLC to the VQA domain. We evaluated its performance on the POPE benchmark [pope], which is designed to probe object existence hallucinations. As presented in Table 2, DLC achieves state-of-the-art results among training-free methods across all three challenging settings: random, popular, and adversarial. This robust performance demonstrates the effectiveness and versatility of our calibration framework, confirming its applicability to more constrained, reasoning-based tasks.
@@
Rand. Pop. Adv. Rand. Pop. Adv. Rand. Pop. Adv.
@@
SHR Evaluation. Following [opera, sid, hadpo], we evaluate on 200 images from the VG dataset using the prompt “Please describe this image in detail” with a maximum of 512 new tokens generated. As illustrated in Figure 3, DLC achieves state-of-the-art performance on hallucination-related metrics at both the sentence and word levels, indicating its effectiveness in mitigating hallucinations driven by strong semantic priors. To characterize generation quality beyond faithfulness, we additionally report a n-gram fluency score (n=2n{=}2) as a proxy for local smoothness and the number of generated sentences per image (SPI) to compare the detailedness of generated texts. Consistent with observations in OPERA [opera], we find that for certain models, DLC may reduce the length of generated descriptions, which is mainly attributable to suppressing superfluous hallucinatory content rather than degrading the core visual description.
@@
Figure 3: SHR evaluation results. Six aspects are analyzed, including the number of sentences per image (SPI), fluency and local smoothness (2-gram), the number of hallucinated sentences per image (HSPI), the number of hallucinated words per image (HWPI), the ratio of hallucinated sentences (HSR), and the ratio of hallucinated words (HWR). Larger radar indicates better performance.
GPT-4o Assisted Evaluation. To comprehensively analyze hallucination and text quality across diverse data distributions, we employ GPT-4o as an advanced evaluation proxy to score Correctness ($C$) and Detailedness ($D$). Our evaluation spans two distinct benchmarks: standard object descriptions on COCO [mscoco] (500 randomly selected images) and the more challenging, out-of-distribution scenarios in LLaVA-Bench-in-the-Wild [llava].The quantitative results, consolidated in Table 3, benchmark DLC against Nucleus Sampling and established decoding strategies including VCD [vcd], OPERA [opera], and SID [sid]. The unified view reveals that DLC achieves remarkable consistency: it outperforms baselines across both benchmarks and all three model architectures. Notably, DLC demonstrates significant gains in correctness over conventional sampling while maintaining comparable descriptive detail. This robustness across datasets suggests that DLC is not merely overfitting to specific object distributions but offers a generalized improvement in grounding. Furthermore, we provide an in-depth qualitative analysis of case studies in Appendix. These cases visually demonstrate how the DLC approach markedly reduces hallucinatory content and enhances descriptive detail under identical image and prompt conditions.
Table 3: Hallucination Evaluation with GPT-4o. We report Correctness and Detailedness across two benchmarks. Grouping by model allows for direct comparison of performance consistency across COCO and LLaVA-Bench-in-the-Wild (L-Bench).